Red, Inc., Yellow Corp., and Blue Company each will pay a dividend of $2.75 next year. The growth rate in dividends for all three companies is 6 percent. The required return for each company's stock is 9.90 percent, 12.70 percent, and 14.70 percent, respectively.

What is the stock price for Red. Inc., Inc.?

What is the stock price for Yellow Corp.?

What is the stock price for Blue Company?
Answer #1

Current price=D1/(Required return-Growth rate)

Current price for:

Red. Inc.=2.75/(0.099-0.06)=$70.51(Approx).

Yellow Corp.=2.75/(0.127-0.06)=$41.04(Approx).

Blue Company=2.75/(0.147-0.06)=$31.61(Approx).